OpenStudy (maleena_1):

The area of a square garden is 200m. How long is the diagonal?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

A square is a rectangle with equal length and width. \(A = LW = L \times L = L^2\) \(\Large A_{square} = L^2\)

OpenStudy (mathstudent55):

\(L^2 = 200\) Take the square root of both sides.

OpenStudy (mathstudent55):

That will give you L, the side of the square. Then you need to find the diagonal.

OpenStudy (maleena_1):

so the square root of 200?

OpenStudy (mathstudent55):

Correct. The side of the square is \(\sqrt{200} \) Now you need the diagonal. Use the Pythagorean theorem.  \(a^2 + b^2 = c^2\)

OpenStudy (mathstudent55):

In this case, both a and b are \(\sqrt{200} \). Plug them in, and solve for c.

OpenStudy (maleena_1):

I got 19.7

OpenStudy (mathstudent55):

You should get a whole number. \((\sqrt{200})^2 + (\sqrt{200})^2 = c^2\) \(200 + 200 = c^2\) \(400 = c^2\) \(c = 20\)

OpenStudy (maleena_1):

ok I see what you did thank you again

OpenStudy (mathstudent55):

You're welcome.